# Amortized Variational Transdimensional Inference

### v1

@@ 5.1 Bayesian misspecified robust variable selection
Figure 2: Quality of VTI approximation for Bayesian misspecified robust variable selection. Top row: Estimated model probabilities $q_{\psi,\phi}(m)$ vs true model probabilities $\pi_{\psi,\phi}(m)$ on the log scale. Bottom row: Cross entropy between individual model estimates $q_{\psi,\phi}(\bm{\theta}_{m}|m)$ and true density $\pi(\bm{\theta}_{m}|m)$ versus true model probability. Colors indicate 10 replicated analyses, each with $|\pazocal{M}|=2^{7}=128$ models. Columns indicate different normalizing flow constructions for a medium and high misspecified model, increasing flow expressivity from left to right. The flow types are a diagonal Gaussian (Diagonal Gaussian MLP), affine masked autoregressive flow composed of 5 layers each with 5 blocks (Affine MAF (5,5)), and a rational quadratic spline masked autoregressive flow composed of 4 layers each with 6 blocks (Spline MAF (4,6)) (see Section˜A.2 for details).
@@
$p\bigl(y_{i}\mid\bm{x}_{i},\bm{\beta},\bm{\gamma}\bigr)$ $=\bm{\mu}(\bm{x}_{i})+(1-\alpha)\,\mathcal{N}\!\Bigl(y_{i};0,\sigma_{1}^{2}\Bigr)\;+\;\alpha\,\mathcal{N}\!\Bigl(y_{i};0,\sigma_{2}^{2}\Bigr),$ (20)
and priors $p(\gamma)\;=\;2^{-p}$ and $p(\bm{\beta})\;=\;\mathcal{N}(0,\,\sigma_{\beta}^{2}\mathbf{I})$. Here, $\alpha$ controls the fraction of outliers, and $(\sigma_{1}^{2},\sigma_{2}^{2})$ encode the variances of in-distribution and outliers, respectively. To complicate the inference problem, two misspecified data-generating processes were used, where a medium-misspecifed DGP sets $\sigma_{1}=2,\sigma_{2}=5$ and a high-misspecified DGP sets $\sigma_{1}=4,\sigma_{2}=4$. This subverts the uni-modality of each $\pi(\bm{\theta}_{m}|m)$ such that each becomes a multi-modal distribution that is difficult to sample and also approximate. Table 2 in Appendix E summarizes the full experiment configuration. Figure 2 shows how increasing complexity of the variational density (left-to-right panels) improves the quality of the approximations of both $\pi(\bm{\theta}_{m}|m)$ (bottom row) and estimated model probabilities (top row), and that the approximation quality of $\pi(\bm{\theta}_{m}|m)$ is higher for higher probability models. Appendix E shows further results for larger $|\pazocal{M}|$, and demonstrates VTI robustness to diffuse priors.
@@ 5.2 Bayesian non-linear directed acyclic graph discovery
We consider a dataset of real-valued observations, denoted by $\bm{X}\in\mathbb{R}^{n\times N_{d}}$, where $n$ is the number of data samples and $N_{d}$ is the number of nodes. Our goal is to perform Bayesian inference over a space of non-linear structural equation models (SEMs) which is isomorophic to a space of directed acyclic graphs (DAGs) and non-linear functions over the active edges. A DAG is represented by a directed adjacency matrix $\bm{A}\in\{0,1\}^{N_{d}\times N_{d}}$, where $A_{ij}=1$ indicates a directed edge from node $i$ to node $j$ and $A_{ij}=0$ otherwise. The acyclicity constraint requires that the directed edges in $\bm{A}$ do not form any directed cycle. In a non-linear SEM, each node $X_{j}$ depends non-linearly on its parents in the form $\bm{X}=f(\bm{X})+\epsilon$, $\epsilon_{j}\sim\mathcal{N}\!\bigl(0,\,\sigma^{2}\bigr)$, where $f:\mathbb{R}^{N_{d}}\mapsto\mathbb{R}^{N_{d}}$ is a nonlinear function possessing an acyclic Jacobian matrix. We follow Bello et al., (2022); Thompson et al., (2025) whereby $f$ is a multi-layer-perceptron (MLP) structured as $f(\bm{X})=(f_{1}(\bm{X}),\dots,f_{N_{d}}(\bm{X}))^{\top}$. We implement $f$ using a single hidden layer, with rectified linear unit (ReLU) activation functions used to model non-linearity where the bias term can be optionally included (see Appendix˜F for details). Acyclicity is enforced by introducing a topological ordering of the $N_{d}$ nodes. By letting $\bm{P}$ be a permutation matrix that reorders nodes into a valid topological order, and defining $\bm{U}$ to be strictly upper-triangular, we can represent any acyclic adjacency matrix as $\bm{A}\;=\;\bm{P}^{\top}\,\bm{U}\,\bm{P}$. Each edge is guaranteed to point from lower-indexed nodes to higher-indexed nodes in the topological order (Bonilla et al., , 2024). Note that this parametrization does not conform to Proposition˜3.1, as the correspondence between $(\bm{P},\bm{U})$ and $\bm{A}$ is not one-to-one. However, we can circumvent this by sampling with a MADE-based discrete distribution (Germain et al., , 2015) $q_{\psi}$ for inference over a very high cardinality model space (see Appendices F.8 and F.7 for details). See Section˜F.4 for details of the simulation study in Figure 3, contrasting VTI with the baseline DAGMA (Bello et al., , 2022).
@@
Figure 3: Simulation study comparing VTI to DAGMA (Bello et al., , 2022) for discovery of a 10-node non-linear DAG visualized using standard metrics (Section˜F.2, left to right, where better is: higher, lower, lower, higher). Bars display mean and standard error over 10 i.i.d. repetitions for each data set size.
@@
Sachs et al., (2005) utilize Bayesian networks to analyze multi-parameter single-cell data for deriving causal influences in cellular signaling networks of human immune cells. Causal interactions are validated by comparing to a domain-scientist “agreed” adjacency matrix representing causality within the data, establishing a baseline for causal prediction accuracy. We use VTI to discover the distribution of non-linear DAGs for these data, comprising $n=7466$ entries over $N_{d}=11$ nodes, and benchmark this against the agreed adjacency. Table 1 shows the strong performance of VTI posterior approximation compared to state of the art methods.
Method F1 SHD Brier AUROC
VTI Non-linear DAG 0.61 15.0 15.0 0.79
DAGMA Non-linear 0.32 25.0 25.0 0.60
Table 1: Comparison of VTI non-linear DAG formulation for parameters on data from Sachs et al., (2005) compared to a baseline DAGMA non-linear (Bello et al., , 2022) approach.
@@ 6 Discussion
We have introduced CoSMIC  normalizing flows as a means to implement amortized variational transdimensional inference (VTI), the approximation of a target density over a transdimensional space with a single variational density. VTI is broadly applicable to a wide class of transdimensional inference problems, and we have shown its utility in representative problems involving misspecified robust variable selection and non-linear DAG discovery. We have presented two approaches for simultaneously optimizing the variational parameters $\psi,\phi$, each with benefits and drawbacks. We have also derived approximation error bounds for our Gaussian surrogate-based approach and established convergence guarantees for the marginal models distribution under convergent optimization steps. The choice of model sampler is dependent on the cardinality of the model space. When $|\pazocal{M}|$ is below the computational memory limit, the Gaussian process surrogate-based sampler is recommended. For high cardinality problems we recommend a neural model sampler for approximate inference on the distribution of model weights using Monte Carlo gradients for SGD optimization.
@@
There are several avenues for future research in CoSMIC  flows and VTI. As the variational objective $\pazocal{L}(\psi,\phi)$ (equation 10) returns a single loss value while taking an expectation over $q_{\psi}(m)$, the VTI approximation quality possesses two notable characteristics. The first is that those models $m\in\pazocal{M}$ estimated to have large posterior model probabilities will contribute most significantly to the loss. Hence the CoSMIC  flow will produce a relatively more accurate (in the KL sense) approximation of such models. Conversely those models with low posterior model probability will have relatively worse variational approximations. This effect is seen in Figure 2 (bottom row), regardless of normalizing flow expressivity. While one might prefer greater accuracy on more dominant models, structured changes to $\pazocal{L}(\psi,\phi)$, e.g. weighted stratifications of models across various partitions, could give greater control over where the quality of the variational approximation should focus. The second characteristic is that when the conditional normalising flow is unable to approximate the conditional target $\pi(\bm{\theta}_{m}\,|\,m)$ well, a smaller loss can be achieved by shrinking the estimated model probability $q_{\psi}(m)$ to zero. This effect is seen in Figure 2 (top row), which lessens as flow expressivity increases. Here the question is how to design the normalizing flow, i.e. the flow context $\bm{\xi}$ and the mapping $C(m)$, to best allocate resources to producing good approximations of those models likely to be of relatively high posterior model probability? E.g., in many transdimensional problems, such as robust misspecified variable section problem (Section 5.1), models adjacent (e.g. by flipping the state of a single $\gamma_{j}$) to models with high posterior model probability are also high model probability candidates.
Future work can also derive convergence rates, which will depend on the choice of optimization algorithm for the flow parameters. In addition, our analysis for the surrogate-based approach is general enough to be extended to a variety of methods for approximating a models’ distribution.
Appendix A Construction and analysis of a CoSMIC normalizing flow
CoSMIC IAF procedure: Draw reference samples

### v2

@@ 5.1 Bayesian misspecified robust variable selection
$p\bigl(y_{i}\mid\bm{x}_{i},\bm{\beta},\bm{\gamma}\bigr)$ $=(1-\alpha)\,\mathcal{N}\!\Bigl(y_{i};\bm{\mu}(\bm{x}_{i}),\sigma_{1}^{2}\Bigr)\;+\;\alpha\,\mathcal{N}\!\Bigl(y_{i};\bm{\mu}(\bm{x}_{i}),\sigma_{2}^{2}\Bigr),$ (16)
and priors $\smash{p(\gamma)\;=\;2^{-p}}$ and $\smash{p(\bm{\beta})\;=\;\mathcal{N}(0,\,\sigma_{\beta}^{2}\mathbf{I})}$. Here, $\alpha$ controls the fraction of outliers, and $(\sigma_{1}^{2},\sigma_{2}^{2})$ encode the variances of in-distribution and outliers, respectively. To complicate the inference problem, two misspecified data-generating processes were used (medium- and high-misspecification) which encourages multi-modality in the approximating posterior $\pi(\bm{\theta}_{m}|m)$.
Table 2 in Appendix F summarizes the full experiment configuration. Figure 2 offers a holistic assessment of inference quality relative to a sampling baseline using RJMCMC, where cross-entropy reduces as flow expressivity increases. It shows two problem settings, mid and high misspecification, and for each setting shows how increasing complexity of the variational density (left-to-right panels) improves the quality of the approximations of both $\pi(\bm{\theta}_{m}|m)$ (bottom row) and estimated model probabilities (top row), and that the approximation quality of $\pi(\bm{\theta}_{m}|m)$ is higher for higher probability models.
Refer to caption  Refer to caption
Figure 2: Quality of VTI approximation for Bayesian misspecified robust variable selection. Outer columns denote medium (left) or high (right) likelihood misspecificaton, inner columns indicate different normalizing flow constructions, increasing flow expressivity from left to right. Flow types are described in Appendix A.2. Top row: Estimated model probabilities $q_{\psi}(m)$ vs true model probabilities $\pi(m)$ on the log scale. Bottom row: Cross entropy between individual model estimates $q_{\phi}(\bm{\theta}_{m}|m)$ and true density $\pi(\bm{\theta}_{m}|m)$ versus true model probability. Colors indicate 10 replicated analyses, each with $|\pazocal{M}|=2^{7}$ models.
Cardinality sweep:
Using the focused prior setup on both the medium and high misspecification level targets, we sweep the cardinality of the model space $|\pazocal{M}|$ from $2^{9}$ to $2^{24}$ and compute the cross entropy $H(\pi,q_{\psi,\phi})$, where samples $(m,\bm{\theta}_{m})\sim\pi$ are obtained via RJMCMC (see Appendix F.3). Figure 3 (left) compares the cross entropy between the three $q_{\psi}(m)$ types discussed in Sections 3.1 and 3.3 in simulated problems of increasing $|\pazocal{M}|$. As expected, $H(\pi,q_{\psi,\phi})$ generally increases with $|\pazocal{M}|$ when the flow architecture is held fixed. The surrogate method (blue bars) performs comparably with the other methods for the smaller model spaces ($|\pazocal{M}|=2^{9}$), whereas the neural density (orange bars) performs consistently as $|\pazocal{M}|$ increases. Figure 3 (right) shows two bivariate plots of selected variables ($\smash{\bm{\theta}_{m}^{(1)},\bm{\theta}_{m}^{(5)})}$ from the posterior inferred using RJMCMC and VTI. This qualitative visual comparison shows how well the CoSMIC flow is able to capture non-trivial model distributions versus the sampling approach (for the full multivariate comparison see Figure 11). Appendix F describes the experiments in further detail and demonstrates VTI robustness to diffuse priors.
Refer to caption Refer to caption  Refer to caption Refer to caption
Figure 3: Left: A simulation study of the robust variable selection example showing the cross entropy (NLL) between RJMCMC samples and an amortized variational transdimensional density using rational quadratic spline CoSMIC flows under a fixed number of iterations (30,000). Each cardinality was run with 10 independently sampled synthetic data sets. Right: Comparison of bivariate plots of variables $\smash{\bm{\theta}_{m}^{(1)},\bm{\theta}_{m}^{(5)}}$ obtained by RJMCMC and VTI for a single $\smash{|\pazocal{M}|=2^{7}}$ problem.
@@ 5.2 Bayesian non-linear directed acyclic graph discovery
We consider a dataset of real-valued observations, denoted by $\bm{X}\in\mathbb{R}^{n\times N_{d}}$, where $n$ is the number of data samples and $N_{d}$ is the number of nodes. Our goal is to perform Bayesian inference over a space of non-linear structural equation models (SEMs) which is isomorphic to a space of directed acyclic graphs (DAGs) and non-linear functions over the active edges. A DAG is represented by a directed adjacency matrix $\mathbf{A}\in\{0,1\}^{N_{d}\times N_{d}}$, where $A_{ij}=1$ indicates a directed edge from node $i$ to node $j$ and $A_{ij}=0$ otherwise. The acyclicity constraint requires that the directed edges in $\mathbf{A}$ do not form any directed cycle. In a non-linear SEM, each node $X_{j}$ depends non-linearly on its parents in the form $\bm{X}=f(\bm{X})+\epsilon$, $\epsilon_{j}\sim\mathcal{N}\!\bigl(0,\,\sigma^{2}\bigr)$, where $f:\mathbb{R}^{N_{d}}\mapsto\mathbb{R}^{N_{d}}$ is a nonlinear function possessing an acyclic Jacobian matrix. We follow [Bello et al., 2022, Thompson et al., 2025] whereby $f$ is a multi-layer-perceptron (MLP) structured as $f(\bm{X})=(f_{1}(\bm{X}),\dots,f_{N_{d}}(\bm{X}))^{\top}$. We implement $f$ using a single hidden layer, with rectified linear unit (ReLU) activation functions used to model non-linearity where the bias term can be optionally included (see Appendix G). By introducing a topological ordering of the $N_{d}$ nodes, we simultaneously enforce acyclicity and a consistent mapping of parameters to each graph. Let $\mathbf{P}$ be a permutation matrix that reorders nodes into a valid topological order and define $\mathbf{U}$ to be strictly upper-triangular. By construction, any acyclic adjacency matrix can be represented as $\mathbf{A}\;=\;\mathbf{P}^{\top}\,\mathbf{U}\,\mathbf{P}$. Each edge is guaranteed to point from lower-indexed nodes to higher-indexed nodes in the topological order [Bonilla et al., 2024]. Note that this parametrization does not conform to Proposition˜D.1, as the correspondence between $(\mathbf{P},\mathbf{U})$ and $\mathbf{A}$ is many-to-one. However, this does not violate the consistent parameter mapping. We use a MADE-based discrete distribution [Germain et al., 2015] for $q_{\psi}$ for inference over a very high cardinality model space (see Appendices G.8 and G.7 for details). The simulation study in Figure 4 contrasts VTI with state-of-the-art Bayesian and non-Bayesian baselines (DiBS/DiBS+ Lorch et al. [2021], JSP-GFlowNets Deleu et al. [2023], and DAGMA [Bello et al., 2022]) with the aim of demonstrating that the performance of the generic VTI approach can be competitive with application-specific approaches, where one would expect the latter to have better performance. Evaluation of each method is depicted using the commonly accepted F1 score, structural Hamming distance (SHD), Brier score, and area under receiver operating curve (AUROC) (see Appendix G.2). A complete description of this study is in Appendix G.4.
@@
Figure 4: Simulation study comparing VTI to DAGMA [Bello et al., 2022], DiBS/DiBS+ Lorch et al. [2021], and JSP-GFlowNets Deleu et al. [2023] for discovery of a 10-node non-linear DAG visualized using standard metrics (Appendix G.2, left to right, where better is: higher, lower, lower, higher). Bars display mean and standard error over nine i.i.d. repetitions for each data set size.
Table 1: Comparison of DAG discovery on flow cytometry data [Sachs et al., 2005]: VTI versus baselines
Method F1 SHD Brier AUROC
VTI non-linear DAG 0.44 23.0 23.0 0.68
DAGMA non-linear 0.32 25.0 25.0 0.60
DiBS+ non-linear 0.22 28.0 17.0 0.54
JSP-GFN non-linear 0.23 54.5 44.0 0.51
@@
Sachs et al. [2005] use Bayesian networks to analyze multi-parameter single-cell data for deriving causal influences in cellular signaling networks of human immune cells. Causal interactions are validated by comparing to a domain-agreed adjacency matrix representing causality within the data, establishing a baseline for causal prediction accuracy. We use VTI to discover the distribution of non-linear DAGs for these data, comprising $n=7466$ entries over $N_{d}=11$ nodes, and benchmark this against the agreed adjacency. Table 1 shows strong performance of VTI compared to state of the art methods. A complete description is in Appendix G.5.
@@ 6 Discussion
We have introduced CoSMIC  normalizing flows as a means to implement amortized variational transdimensional inference (VTI), the approximation of a target density over a transdimensional space with a single variational density. VTI is broadly applicable to a wide class of transdimensional inference problems. Although the specification of a CoSMIC flow requires augmenting all model dimensions to $d_{\mathrm{max}}$}, VTI is not sensitive to these added dimensions during training and inference due to the construction of the auxiliary variable transforms. We presented two approaches for simultaneously optimizing the variational parameters $\psi,\phi$. The Gaussian surrogate-based approach benefits from our derivation of the approximation error bounds and established convergence guarantees for the marginal models distribution under convergent optimization steps. The two approaches that use Monte Carlo gradient estimation for SGD optimization benefit from recent advances in neural architectures and neural approximation of very large model spaces. The choice of model sampler is dependent on both the cardinality of the model space and the structure of the problem. When $|\pazocal{M}|$ is small, the Gaussian process surrogate-based sampler or the categorical sampler using Monte Carlo gradients are both appropriate, although in practice it is usually safe to default to the latter approach. For high cardinality problems, we recommend a neural model sampler for approximate inference on the distribution of model weights.
@@
The quality of the VTI approximation possesses two notable characteristics. Firstly, those models $m\in\pazocal{M}$ estimated to have large posterior model probabilities will contribute most significantly to the loss. Hence the CoSMIC  flow will produce a relatively more accurate (in the KL sense) approximation of such models, compared to models with low probabilities. This effect is seen in Figure 2 (bottom row). While one might prefer greater accuracy on more dominant models, structured changes to $\pazocal{L}(\psi,\phi)$ could give greater control over where the quality of the variational approximation should focus. The second characteristic is that when the normalizing flow is unable to approximate the conditional target $\pi(\bm{\theta}_{m}\,|\,m)$ well, a smaller loss can be achieved by shrinking the estimated model probability $q_{\psi}(m)$ to zero. This effect is seen in Figure 2 (top row), which lessens as flow expressivity increases. Here the question is how to design the normalizing flow, i.e. the flow context $\bm{\xi}$ and the mapping $C(m)$, to best allocate resources to produce good approximations of models likely to be of relatively high posterior model probability. In many transdimensional problems, two models could be considered adjacent by structural similarity (e.g. in variable selection where they differ by one included covariate) and so may have similar posterior model probabilities. This could be achieved by e.g. extending the architecture of the context encoder (Appendix A.2) to capture similarities between models that generalize over the model space, and learning structural similarity within the surrogate-based model sampler itself via a reward-based criterion.
The left-align permutation used in Proposition 2.2 raises the question of whether the alignment of variables across models in the normalizing flow is important for computational efficiency. More broadly: (a) is there shared information between models? And, if so, (b) would careful manual construction of the flow improve exploitation of this versus allowing the optimization to determine it agnostically? (a) is answered by the robust variable selection example in Figures 3 (right) and 11 where high probability mass regions for each model do not overlap and thus there there is no such shared information. To answer (b) would be an avenue for future research. Future work could also derive convergence rates, which will depend on the choice of optimization algorithm for the flow parameters, and extension of the CoSMIC architecture to coupling flows, in applications outside variational inference. In addition, our analysis for the surrogate-based approach is general enough to be extended to a variety of methods for approximating a distribution over models. Finally, in extending VI to the transdimensional setting we inherit the same strengths and weaknesses of single-model VI, including the challenges of mode-collapse. Users would need to take the same steps to manage it as in the standard setting. Mode collapse in the model distribution is mitigated by the exploration versus exploitation strategies discussed in Section 3.
Acknowledgments
This research was made possible by funding support from the CSIRO Machine Learning and Artificial Intelligence Future Science Platform. The authors would also like to thank Ryan Thompson for critical advice in the construction of the non-linear DAG example, and Edwin Bonilla for advice on Monte Carlo gradient estimation methods. SAS is supported by the Australian Research Council.
Appendix A Implementation of CoSMIC normalizing flows
A.1 Inverse autoregressive flow sampling procedure
Draw reference samples
@@ A.1 Inverse autoregressive flow sampling procedure
A.2 Experimental CoSMIC transform compositions
The experiments use the below compositions of transforms as inverse autoregressive flows $T_{\phi}(\bm{z}\mid m)$ where $\bm{z}$ are the inputs from the reference distribution and $m$ is the context input. All compositions except for the diagonal Gaussian are assumed to have the strict left-align permutations discussed in Appendix A. The term “block” is defined in Appendix A.3.
Context encoder:
Model-specific reverse-permutation:
Flow compositions commonly include reverse permutations to ensure expressibility of an autoregressive-NN-based flow is (approximately) the same for all coordinates. Denoting the generic reverse permutation for all coordinates as $P^{\operatorname{rev}}$, we assume the strict left-right permutation $P_{m}$ (as per Appendix A) has been applied, and hence define the left-most $d_{m}$-coordinate reverse permutation $P_{<d_{m}}^{\operatorname{rev}}$.
Affine(5,5):
The learned component is the affine masked autoregressive transform [Papamakarios et al., 2017], denoted here as $T_{\phi_{k}}^{\operatorname{Affine}}$ for transforms $k=1,\dots,5$, each having 5 blocks. We set $T_{\phi}:=T_{\phi_{5}}^{\operatorname{Affine}}\circ P_{<d_{m}}^{\operatorname{rev}}\circ\dots\circ P_{<d_{m}}^{\operatorname{rev}}\circ T_{\phi_{1}}^{\operatorname{Affine}}$.
Spline(4,6):
The learned component is the rational quadratic spline masked autoregressive flow architecture [Durkan et al., 2019], denoted here as $T_{\phi_{k}}^{\operatorname{RQ-Spline}}$. Each $T_{\phi_{k}}^{\operatorname{RQ-Spline}}$ has 6 blocks. Additionally, we define a fixed global affine transform $T_{\mu_{g},\sigma_{g}}$ that is not dependent on inputs nor context and hence has only two learnable parameters: scale $\mu_{g}$ and shift $\sigma_{g}$. We set $T_{\phi}:=T_{\mu_{g},\sigma_{g}}\circ T_{\phi_{5}}^{\operatorname{RQ-Spline}}\circ P_{<d_{m}}^{\operatorname{rev}}\circ\dots\circ P_{<d_{m}}^{\operatorname{rev}}\circ T_{\phi_{1}}^{\operatorname{RQ-Spline}}$.
A.3 Autoregressive flow definitions
We use the residual variant of the Masked Autoencoder for Distribution Estimation (MADE) [Germain et al., 2015], implemented in PyTorch by [Durkan et al., 2020]. Each block maintains the autoregressive property by assigning degrees $\{1,\dots,d\}$ to inputs and propagating them forward.
Appendix B Analysis of a CoSMIC normalizing flow
Proof of Lemma˜2.1.
Let $I(m)=\{\,i\in\{1,\dots,{d_{\mathrm{max}}}\}:A_{i}(m)=1\,\}$, and $I^{c}$ be the complement. The result holds from equation 7 as, for all coordinates $i\in I^{c}(m)$, $\bm{u}_{\setminus m}^{(i)}=\tau_{\bm{\rho}_{i}^{C}}(\bm{z}_{\setminus m}^{(i)})=\tau_{\bm{\rho}^{\mathrm{Id}}}(\bm{z}_{\setminus m}^{(i)})=\bm{z}_{\setminus m}^{(i)}.$